ψ=squar 2/L sin (nPie x/2) find the value of x

1
Expert's answer
2022-01-10T13:48:29-0500

"\psi=\sqrt{\frac 2L}\sin(\frac {n\pi x}2),"

"-\pi≤\frac{n\pi x}2≤\pi,"

"-2\u2264nx\u22642,"

"-\frac 2n≤x≤\frac 2n,~n=1,2,3,..."